OpenStudy (bbb911):

___3x - 1___ - ____x - 3____ 2x x subtract and simplify factor will give medal

OpenStudy (anonymous):

holy moly lets see if we can read it

OpenStudy (anonymous):

\[\frac{3x-1}{2x}-\frac{x-3}{x}\]?

OpenStudy (bbb911):

sorry the under line is like the dividing yes correct

OpenStudy (anonymous):

then the common denominator will be \(2x\) and you need to compute \[\frac{3x-1-2(x-3)}{2x}\]